Average cash $56,250
Average accounts payable $540,000
Average accounts receivable $1,350,000
Average inventories $675,000
Average cash sales $4,500,000
Average credit sales $13,500,000
Average cost of goods sold $8,100,000
Average number of days per year 365 days
Inventory conversion period 30.42 days
Payables deferral period 1. days A. 24 B. 24.33 C. 10.95.?
Receivables conversion period 2. days A 36.5 B. 36 C. 27.38 ?
Cash conversion cycle 3. days A. 54.75 B. 91.25 C. 42.59 ?
Answer #1

Payables deferral period = 365 * Average account payables / Average cost of goods sold

= 365 * 540000 / 8100000 = 24.33 days

Receivables conversion period = 365 * Average account receivables / Average credit sales

= 365 * 1350000 / 13500000

= 36.50 days

Inventory conversion period (given) = 30. 42 days

Cash conversion cycle = Inventory period + Receivables period - Payables deferral period

= 30.42 + 36.50 - 24.33

= 42.59 days